OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1? infinitely many solutions no solution one solution

OpenStudy (xapproachesinfinity):

x+y=1 and y=-x+1 those are the same equation

OpenStudy (xapproachesinfinity):

how many points that lie on that line satisfying x+y=1

OpenStudy (xapproachesinfinity):

what do you think?

OpenStudy (xapproachesinfinity):

how many points would satisfy x+y=1 just use some logic!!

OpenStudy (anonymous):

one solution

OpenStudy (xapproachesinfinity):

no! there are so many points that give us x+y=1

OpenStudy (anonymous):

infinitely many solutions then

OpenStudy (xapproachesinfinity):

one solution is when we have two distinct line intersecting on one points that point is the solution to the system determined by the equations of those two line

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

but i don't think you got this yet

OpenStudy (xapproachesinfinity):

what if two lines are parallel to each other how many solutions will there be?

OpenStudy (anonymous):

many

OpenStudy (xapproachesinfinity):

hmm... first you need to know that there is a solution only when there is intersection ====== many solution only in one case where two lines coincide (meaning the same line )

OpenStudy (xapproachesinfinity):

so when two line are parallel there is no solution